# PyRadiomics-cuda: a GPU-accelerated 3D features extraction from medical images within PyRadiomics

### v1

@@ 4 Conclusions
The initial motivation for this work was long time of feature extraction needed for training AI model in the xLUNGS project [6] dealing with approximately 40000 lungs CT (Computer Tomography) scans [4], perhaps the world’s largest chest X-ray database. Thanks to PyRadiomics-cuda the 3D features extraction time was significantly reduced which led to crucial savings in expensive infrastructure utilization costs. PyRadiomics-cuda proved to deliver the same quality output as original PyRadiomics, but in a fraction of the time not requiring any changes to the existing workflows.
5 Competing interests
No competing interest is declared.
6 Author contributions statement
J.L. and P.T. implemented the tool, created the public repository, testing framework and performed the experiments. J.L., S.Z. and K.K. worked on the text, data analysis and presentation.
7 Acknowledgments
This research was carried out with the support of the High Performance Computing Center at Faculty of Mathematics and Information Science Warsaw University of Technology.

### v2

@@ 4 Conclusions
The initial motivation for this work was long time of feature extraction needed for training AI model in the xLUNGS project [Luckner2025] dealing with approximately 40000 lungs CT (Computer Tomography) scans [kolodko_polish_2025], perhaps the world’s largest chest X-ray database. PyRadiomics-cuda significantly reduced 3D feature extraction time, leading to substantial savings in expensive infrastructure utilization, while delivering output with identical quality to the original PyRadiomics without requiring any changes to existing workflows.
We showed that PyRadiomics-cuda tackles the scalability and performance limitations of radiomics feature extraction, a core component of many biomedical imaging pipelines. By leveraging GPU acceleration while preserving compatibility with a widely adopted biomedical software framework, it bridges high-performance computing techniques and practical biomedical data analysis. The proposed solution enables efficient large-scale studies, supports reproducible research through open-source release, and provides a foundation for future extensions to distributed and multi-modal biomedical workflows.
PyRadiomics-cuda is implemented in Python and C/CUDA and as PyRadiomics is freely available under the BSD license at GitHub [MiswutPyradiomicsCUDA2026].
Funding.
This research was carried out with the support of the High Performance Computing Center at Faculty of Math. and Inform. Science and POB Cybersecurity and data analysis of Warsaw Univ. of Techn. within the Excellence Initiative: Research University (IDUB) program.